# A Survey of NL2SQL with Large Language Models: Where are we, and where are we going?

### v4

@@ IV-C2 Retrieval-based Methods
Retrieval-based methods improve the effectiveness of acquiring additional information but increase computing costs. Moreover, current research mostly relies on domain-specific text, with limited use of structured knowledge. Thus, integrating diverse information sources could further enhance nl2sql performance, especially for domain-specific databases.
V NL2SQL Translation Methods
In this section, we elaborate on nl2sql translation methods using language models. As shown in Figure 5, we will detail their encoding (Section V-A), decoding (Section V-B), and task-specific prompt strategies (Section V-C). Moreover, we will discuss how the intermediate representation can benefit the nl2sql translation process (Section V-D).
@@ V-A Encoding Strategy
In the nl2sql task, encoding refers to transforming nl and database schema into a structured format suitable for language model processing. This step is essential for converting unstructured data into a form usable for sql generation, capturing the nl’s semantics and the schema’s structure to help the model map user intent to appropriate sql. As shown in Figure 6, primary encoding strategies include 1) Sequential Encoding, 2) Graph-based Encoding, and 3) Separate Encoding.
@@ V-A1 Sequential Encoding Strategy
Sequential encoding is a strategy in nl2sql models where both the nl and the database schema are treated as token sequences. The key idea is to linearize the input data, enabling sequence-based models to effectively capture semantic and syntactic information, as shown in Figure 6(a). For example, models like T5 [42] are used to encode nl and database schema sequentially in works [63, 73] BRIDGE [79] improves the alignment between the nl and database schema by representing both as a tagged sequence and inserting matched database cell values (called anchor texts) next to corresponding fields. Similarly, N-best List Rerankers [71] appends database content to column names. RESDSQL [7] uses a ranking-enhanced encoder to sort and filter schema items, prioritizing the most relevant ones and reducing schema linking complexity. CatSQL [67] leverages the pre-trained GraPPa encoding network [37] to concatenate nl, schema, and additional information into an input sequence.
Although LLM-based nl2sql methods don’t explicitly define an input encoding strategy, most rely on processing input sequences using the self-attention mechanism. This mechanism functions similarly to an encoding process, predicting the next word based on preceding ones in the sequence.
While sequential encoding is simple and intuitive, it may struggle to capture the complex relationships between the database schema and the nl query, potentially limiting the model’s ability to understand and generate more complex sql.
@@ V-A2 Graph-based Encoding Strategy
Graph-based encoding in nl2sql models represents both nl and database schema as interconnected graphs, leveraging the relational structure of databases and inter-dependencies in the input data, as shown in Figure 6(b) [76, 77, 37, 78, 75, 72, 70, 68, 43]. Unlike sequential encoding, this approach preserves the schema’s topology, offering richer context for each element and enhancing the model’s ability to produce accurate sql queries.
RAT-SQL [76] introduces a relation-aware self-attention mechanism, explicitly using relational information in a graph structure to jointly encode the question and schema, enhancing the model’s understanding of structural information. S2SQL [75] injects syntactic structure information at the encoding stage using the ELECTRA [103] model, enhancing semantic understanding. G3[69] uses the LGESQL[104] encoder and Graph Attention Network (GAT) [105] to capture multi-source heterogeneous information. SQLformer [68] introduces learnable embeddings for tables and columns to select the most relevant schema elements.
Graph-based encoding effectively captures complex relationships between input data, enabling the model to generate more accurate sql queries with multiple relationships and conditions. However, it requires more sophisticated algorithms for graph construction and processing, and its full potential may only be realized with large amounts of training data, making it less suitable for data-limited scenarios.
@@ V-A3 Separate Encoding Strategy
Separate encoding strategy is an approach in nl2sql models where different parts of the nl (such as clauses and conditions) are encoded separately and then combined to form the final sql, as shown in Figure 6(c).
Early models, such as SQLNet [106] and Seq2SQL [107], used separate encoding for the nl and database schema due to the format mismatch between the two. However, this approach hindered schema linking and is now less commonly used. Despite this, separate encoding allows for tailored handling of different data types, leveraging the strengths of various encoding techniques. TKK [74] employs task decomposition and multi-task learning strategies by breaking down the complex nl2sql task into subtasks and progressively integrating knowledge. Similarly, SC-Prompt [1] divides text encoding into two stages: structure and content, each encoded separately.
While the separate encoding strategy requires multiple processing of input data, which may extend the training and inference time of the model, it allows for more refined handling and understanding of different aspects of queries. This provides the model with the flexibility to handle various query tasks, thereby enhancing overall performance.
@@ V-B Decoding Strategy
Decoding is a crucial step in nl2sql translation, transforming encoder-generated representations into sql. An effective decoding strategy ensures that the generated sql are not only syntactically correct but also semantically aligned with the nl queries, while optimizing sql execution efficiency. Figure 7 introduces several key decoding strategies, as discussed below.
@@ V-B1 Greedy Search-based Decoding Strategy
The greedy search-based decoding strategy is simple and fast, selecting the token with the highest probability at each decoding step. This strategy builds the final output sequence by continuously choosing the locally optimal solution, as shown in Figure 7(a).
Since GPT models (e.g., GPT-4) default to greedy search-based decoding, many nl2sql solutions using GPT fall into this category. DTS-SQL [55], based on DeepSeek LLM [108], uses the same approach. Early models like SQLNet [106] and Seq2SQL [107] also rely on greedy search for sql generation.
Greedy search is popular due to its fast decoding speed and simplicity. However, it only considers the locally optimal solution, potentially overlooking long-term dependencies and global optimization. This can lead to suboptimal sql queries, particularly for complex queries, where early mistakes may propagate and accumulate throughout the decoding process.
@@ V-B2 Beam Search-based Decoding Strategy
This strategy that explores a larger search space based on beam search, leading to potentially better results. Instead of selecting only one optimal token at each step, beam search retains multiple candidate sequences (“beams”) and expands the top-$k$ tokens with the highest probabilities until a complete output sequence is generated, as shown in Figure 7(b).
Given its advantages, several nl2sql models employ beam search [78, 67, 7]. RAT-SQL [76] combines relation-aware graph structure encoding with beam search to generate multiple sql candidates, reranking them based on graph structure information. Unlike RAT-SQL, EditSQL [96] uses beam search alongside dialogue history to generate and refine candidate sql queries. SmBoP [77] employs a semi-autoregressive bottom-up decoding approach, improving efficiency by parallelizing sub-tree construction and scoring, with logarithmic time complexity. ZeroNL2SQL [44] retains the top-$k$ hypotheses during the sql sketch generation stage, which are then refined for query and predicate calibration.
Compared to greedy search, beam search retains multiple candidate sequences, which increases memory and computational demands, slowing decoding speed. However, by considering multiple candidates at each step, beam search explores a wider search space, making it more likely to generate accurate and complex sql queries.
@@ V-B3 Constraint-aware Incremental Decoding Strategy
The constraint-aware incremental decoding strategy, introduced by PICARD [73] (Parsing Incrementally for Constrained Auto-Regressive Decoding), is designed for nl2sql tasks. This strategy aims to ensure the generation of syntactically correct sql queries by incorporating constraints during the decoding process, as shown in Figure 7(c).
Roughly speaking, nl2sql models using this strategy generate sql queries incrementally, applying sql grammar rules at each step. As the model predicts the next token, it also checks the syntactic correctness of the partial query against a set of constraints derived from sql grammar rules. This ensures that each token adheres to the required syntax, reducing the chances of generating invalid or incomplete sql queries and improving overall translation accuracy and reliability. PICARD [73] ensures that each token added adheres to the correct sql syntax, significantly reducing errors and incomplete queries. Many models [74, 72, 73, 70, 71, 43] have demonstrated better performance by using PICARD’s decoding strategy.
While this strategy requires more computational resources and processing time due to incremental decoding and constraint application, it effectively ensures syntactic correctness, making it ideal for generating structurally complex sql queries while minimizing errors.
V-B4 Other Decoding Strategies
Beyond the three common decoding strategies, some models introduce specialized approaches to improve decoding accuracy. For example, BRIDGE [79] introduces Schema-Consistency Guided Decoding, which ensures that generated sql queries align with the database schema. It does this by continuously verifying the match between the query and schema during decoding and adjusting the decoding path accordingly based on the results.
@@ V-C Task-specific Prompt Strategy
In the era of LLMs, prompt engineering has become a powerful method for harnessing LLM capabilities across diverse tasks [109]. In nl2sql, task-specific prompts are crafted to guide LLMs in optimizing nl2sql translation, enhancing the accuracy of translating complex nl queries into precise sql queries. Broadly speaking, there are two main types of task-specific prompt strategies: 1) Chain-of-Thought prompting, and 2) Decomposition Strategy.
@@ V-C1 Chain-of-Thought (CoT) Prompting
The CoT prompting [110], known for its effectiveness, showcases the LLM’s reasoning process, improving both the accuracy and interpretability of the generated results. In nl2sql tasks, CoT enhances model performance and ensures that the generated sql statements are more aligned with human expectations [111]. For example, CHESS [54] transforms nl into sql statements through a streamlined pipeline that utilizes LLMs and CoT. This process includes entity and context retrieval, schema selection, sql generation, and revision.
In addition, the integration of CoT with other techniques can enhance the performance of nl2sql models. These techniques include in-context learning [64, 58], logical synthesis [56], calibration with hints [62, 69] and multi-agent system [66]. Specifically, in-context learning and logical synthesis enrich CoT by embedding a deeper linguistic understanding, enabling precise semantic mapping to sql constructs [64, 58]. Calibration with hints fine-tunes model responses, aligning them closely with nl nuances for accurate intent translation [62, 69]. Furthermore, integrating the multi-agent framework with CoT fosters a collaborative approach, with specialized agents handling tasks like schema linking and sql generation, which speeds up reasoning and enhances adaptability [66].
Overall, these techniques create a more robust nl2sql framework, offering better precision and reliability in translating complex nl queries into accurate sql statements.
@@ V-C2 Decomposition Strategy
The decomposition strategy divides the nl2sql task into sequential subtasks, allowing each sub-module to concentrate on a specific generation step, thereby enhancing accuracy, quality, and interpretability.
Different approaches vary in subtask decomposition granularity [55, 74, 69, 61, 66]. For example, TKK [74] applies finer-grained decomposition by breaking down nl2sql parsing into subtasks like mapping nl to SELECT, FROM, and WHERE clauses. This approach helps the model concentrate on each clause, enhancing understanding of the problem, schema, and sql alignment. Similar strategies are used in G3[69] and DEA-SQL [61]. Moreover, decomposition also reduces model complexity. For example, MAC-SQL [66] introduces a Decomposer agent to split the user’s query into subproblems, making sql generation for each part more manageable.
In general, the decomposition strategy divides the nl2sql translation task into multiple subtasks, enabling each sub-module to focus on enhancing its specific output. However, this approach also raises computational costs, making model training and deployment more complex and resource-intensive.
@@
V-D Intermediate Representation for NL2SQL Translation
The nl2sql task is challenging due to the complexity and ambiguity of nl queries, coupled with the syntax-constrained nature of sql. To simplify this process, researchers have developed a grammar-free intermediate representation (IR) to bridge the “free-form” nl question and the “constrained and formal” sql. This IR provides a structured yet flexible format, capturing the essential components and relationships within an nl query without the strict syntax requirements of sql. Figure 8 shows two types of IR strategies, discussed below.
@@ V-D1 SQL-like syntax language
As shown in Figure 8(a), SQL-like syntax language is a simplified sql-like structure. Early approaches used information retrieval techniques to map the original question and schema data into this syntax [93, 112]. Subsequent research efforts have focused on consolidating or eliminating partial clauses or operations in sql queries to simplify SQL-like syntax language [95, 94]. For example, EditSQL [96] adds WHERE and HAVING conditions. Recent research has focused on simplifying syntax languages to improve parsing efficiency [113]. NatSQL (Natural SQL) [97], a widely used SQL-like syntax language, eliminates uncommon sql operators and keywords, streamlining schema linking by minimizing necessary schema items. Combined with PLMs, NatSQL has achieved strong results on various benchmarks [7, 63]. Additionally, Question Decomposition Meaning Representation [114, 98] breaks down the original nl into atomic questions, each serving as an intermediate representation that translates into small formal operations like selecting entities, retrieving attributes, or aggregating data.
@@ V-D2 SQL-like sketch structure
Early works applied fixed sketch rules and neural networks to map the nl into SQL-like sketch structure [115, 94]. For example, SyntaxSQLNet [94] uses a syntax tree and a corresponding decoder, dividing the decoding into nine sub-modules that separately predict operators, keywords, and entities before combining them to generate the final sql. In recent years, the development of language models has allowed researchers to design more elaborate SQL-like sketch structures for parsing [1, 67, 44, 56]. CatSQL [67] constructs a more general template sketch with slots serving as initial placeholders. Its base model focuses on the parsing of nl to fill these placeholders, consequently decreasing the computational cost. Moreover, several recent works cover both SQL-like syntax language and SQL-like sketch transition methods. For instance, RESDSQL [7] introduces a rank-enhanced encoding and skeleton-aware decoding framework. During the decoding phase, its decoder initially generates the sql skeleton and then transforms it into the actual sql query. When combined with NatSQL, RESDSQL demonstrates the ability further to enhance the quality of sql query generation.
@@
VI Post-Processing Strategies for NL2SQL
After the nl2sql model generates the sql, post-processing can refine it to better meet user expectations. This step involves leveraging additional information or models to enhance the sql, with a focus on sql correction, ensuring output consistency, and execution-guided checking.

### v5

@@ IV-C2 Retrieval-based Methods
Retrieval-based methods improve the effectiveness of acquiring additional information but increase computing costs. Moreover, current research mostly relies on domain-specific text, with limited use of structured knowledge. Thus, integrating diverse information sources could further enhance Text-to-SQL performance, especially for domain-specific databases.
V Text-to-SQL Translation Methods
In this section, we elaborate on Text-to-SQL translation methods using language models. As shown in Figure 5, we will detail their encoding (Section V-A), decoding (Section V-B), and task-specific prompt strategies (Section V-C). Moreover, we will discuss how the intermediate representation can benefit the Text-to-SQL translation process (Section V-D).
@@ V-A Encoding Strategy
In the Text-to-SQL task, encoding refers to transforming nl and database schema into a structured format suitable for language model processing. This step is essential for converting unstructured data into a form usable for sql generation, capturing the nl’s semantics and the schema’s structure to help the model map user intent to appropriate sql. As shown in Figure 6, primary encoding strategies include 1) Sequential Encoding, 2) Graph-based Encoding, and 3) Separate Encoding.
@@ V-A1 Sequential Encoding Strategy
Sequential encoding is a strategy in Text-to-SQL models where both the nl and the database schema are treated as token sequences. As shown in Figure 6(a), the model processes the entire input as a linear token sequence using standard Transformer-based architectures.
Models like T5 [42] are used to encode nl and database schema sequentially in works [68, 77]. BRIDGE [83] improves the alignment between the nl and database schema by representing both as a tagged sequence and inserting matched database cell values (called anchor texts) next to corresponding fields. RESDSQL [7] uses a ranking-enhanced encoder to sort and filter schema items, prioritizing the most relevant ones and reducing schema linking complexity. Although LLM-based Text-to-SQL systems often do not explicitly define an input encoding strategy, they typically implicitly adopt a sequential form by concatenating queries and schema components. These models leverage self-attention to model dependencies across the entire sequence.
While this allows for flexible contextualization, such approaches may struggle to capture complex relational structures, limiting their performance on deeply nested sql queries.
@@ V-A2 Graph-based Encoding Strategy
Graph-based encoding in Text-to-SQL models represents both nl and database schema as interconnected graphs, leveraging the relational structure of databases and inter-dependencies in the input data, as shown in Figure 6(b). Unlike sequential encoding, this approach preserves the schema’s topology, offering richer context for each element and enhancing the model’s ability to produce accurate sql queries  [80, 81, 79, 76, 74, 72, 43].
RAT-SQL [80] introduces a relation-aware self-attention mechanism, explicitly using relational information in a graph structure to jointly encode the question and schema, enhancing the model’s understanding of structural information. S2SQL [79] injects syntactic structure information at the encoding stage using the ELECTRA [106] model, enhancing semantic understanding. G3[73] uses the LGESQL[107] encoder and Graph Attention Network (GAT) [108] to capture multi-source heterogeneous information. Graphix-T5 [43] adds graph-aware layers, incorporating structural information directly into the encoding process, significantly improving sql query generation across multiple benchmarks.
However, this strategy typically involves more intricate graph construction and processing algorithms. It also tends to rely on large-scale training data to achieve optimal performance, which limits its applicability in low-resource scenarios.
@@ V-A3 Separate Encoding Strategy
The separate encoding strategy in Text-to-SQL refers to independently encoding different parts of the input (typically nl and the db schema) rather than combining them into a single sequence. This strategy has evolved significantly over time and can be broadly categorized into traditional and modern forms.
In traditional separate encoding, early models like SQLNet [109] and Seq2SQL [110], processed the nl and database schema separately due to format mismatches. However, this lack of interaction between the two components hindered effective schema linking, leading to limited performance and making this approach less common in recent research. Modern separate encoding strategies, as illustrated in Figure 6(c), focus on modular representation learning by decomposing the Text-to-SQL task into subtasks and encoding different aspects separately. TKK [78] employs task decomposition and multi-task learning strategies by breaking down the complex Text-to-SQL task into subtasks and progressively integrating knowledge. Similarly, SC-Prompt [1] divides text encoding into two stages: structure and content, each encoded separately.
While separate encoding may increase computational overhead due to multiple processing steps, it allows for more refined understanding of different aspects of queries. This modularity provides the model with greater flexibility to handle various query tasks, thereby enhancing overall performance.
@@ V-B Decoding Strategy
Decoding is a crucial step in Text-to-SQL translation, transforming encoder-generated representations into sql. An effective decoding strategy ensures that the generated sql are not only syntactically correct but also semantically aligned with the nl queries, while optimizing sql execution efficiency. Figure 7 introduces several key decoding strategies.
@@ V-B1 Greedy Search-based Decoding Strategy
The greedy search-based decoding strategy is a simple and efficient method that selects the token with the highest probability at each decoding step. As illustrated in Figure 7(a), it constructs the output sequence by making a series of locally optimal choices, without considering future possibilities.
Since GPT models (e.g., GPT-4) default to greedy search-based decoding, many Text-to-SQL solutions using GPT fall into this category. DTS-SQL [60], based on DeepSeek [111], uses the same approach. Early models like SQLNet [109] and Seq2SQL [110] also rely on greedy search for sql generation.
Despite its efficiency, greedy search has notable limitations. By focusing only on immediate token probabilities, it fails to account for long-term dependencies or global sequence coherence. As a result, errors made early in the decoding process may propagate and lead to suboptimal or incorrect SQL queries, particularly for complex or multi-step questions.
@@ V-B2 Beam Search-based Decoding Strategy
Beam search is a widely used decoding strategy that explores a broader search space compared to greedy decoding, often leading to higher-quality results. Instead of selecting only the top token at each step, it retains a fixed number of top-ranked partial sequences (known as beams) and expands each by considering the top-$k$ most probable next tokens, as illustrated in Figure 7(b).
Given its advantages, several Text-to-SQL models employ beam search [82, 71, 7]. RAT-SQL [80] combines relation-aware graph structure encoding with beam search to generate multiple sql candidates, reranking them based on graph structure information. Unlike RAT-SQL, EditSQL [99] uses beam search alongside dialogue history to generate and refine candidate sql queries. SmBoP [81] employs a semi-autoregressive bottom-up decoding approach, improving efficiency by parallelizing sub-tree construction and scoring, with logarithmic time complexity. ZeroNL2SQL [44] retains the top-$k$ hypotheses during the sql sketch generation stage, which are then refined for query and predicate calibration.
Compared to greedy decoding, beam search improves the ability to generate syntactically and semantically valid sql queries, especially in complex scenarios, by considering multiple hypotheses at each step. However, this benefit comes at the cost of increased computational complexity and memory usage, potentially slowing down the decoding process.
@@ V-B3 Constraint-aware Incremental Decoding Strategy
Constraint-aware incremental decoding strategies aim to ensure the structural and syntactic validity of sql queries by applying explicit constraints during the decoding process. As shown in Figure 7(c), these strategies incrementally generate sql while enforcing sql grammar constraints at each step.
A representative implementation is PICARD [77] (Parsing Incrementally for Constrained Auto-Regressive Decoding), which integrates sql grammar constraints into the decoding loop. It verifies the syntactic validity of the partially generated query at every step, ensuring that each token adheres to the sql grammar. This significantly reduces the generation of invalid or incomplete queries. Many models [78, 76, 77, 74, 75, 43] have adopted this paradigm to improve performance.
In addition to grammar-level constraints, some models incorporate schema-level constraints during decoding. BRIDGE [83] introduces Schema-Consistency Guided Decoding, which enforces alignment between the generated sql query and the underlying database schema by verifying their consistency and adjusting the decoding path accordingly.
While this strategy introduces additional computational overhead due to per-token constraint evaluation, it offers strong guarantees of syntactic correctness and is particularly effective for generating structurally complex sql queries.
@@ V-C Task-specific Prompt Strategy
In the era of LLMs, prompt engineering has become a powerful method for harnessing LLM capabilities across diverse tasks. In Text-to-SQL, task-specific prompts are crafted to guide LLMs in optimizing Text-to-SQL translation, enhancing the accuracy of translating complex nl queries into precise sql queries. Broadly speaking, there are two main types of task-specific prompt strategies: 1) Chain-of-Thought prompting, and 2) Decomposition Strategy.
@@ V-C1 Chain-of-Thought (CoT) Prompting
The CoT prompting [112], known for its effectiveness, showcases the LLM’s reasoning process, improving both the accuracy and interpretability of the generated results. In Text-to-SQL, CoT enhances model performance and ensures that the generated sql statements are more aligned with human expectations [113].
CHESS [59] transforms nl into sql statements through a streamlined pipeline that utilizes LLMs and CoT. This process includes entity and context retrieval, schema selection, sql generation, and revision. In addition, the integration of CoT with other techniques can enhance the performance of Text-to-SQL models. These techniques include in-context learning [69, 63], logical synthesis [61], calibration with hints [67, 73] and multi-agent system [54]. Specifically, in-context learning and logical synthesis enrich CoT by embedding a deeper linguistic understanding, enabling precise semantic mapping to sql constructs [69, 63]. Calibration with hints fine-tunes model responses, aligning them closely with nl nuances for accurate intent translation [67, 73]. Furthermore, integrating the multi-agent framework with CoT fosters a collaborative approach, with specialized agents handling tasks like schema linking and sql generation, which speeds up reasoning and enhances adaptability [54].
Overall, these techniques create a more robust Text-to-SQL framework, offering better precision and reliability in translating complex nl queries into accurate sql statements. However, CoT prompting may introduce longer reasoning chains and latency, and its effectiveness can be sensitive to prompt design and task complexity.
@@ V-C2 Decomposition Strategy
The decomposition strategy divides the Text-to-SQL task into sequential subtasks, allowing each sub-module to concentrate on a specific generation step, thereby enhancing accuracy, quality, and interpretability.
Different approaches vary in subtask decomposition granularity [60, 78, 73, 66, 54]. TKK [78] applies finer-grained decomposition by breaking down Text-to-SQL parsing into subtasks like mapping nl to SELECT, FROM, and WHERE clauses. This approach helps the model concentrate on each clause, enhancing understanding of the problem, schema, and sql alignment. Similar strategies are used in G3[73] and DEA-SQL [66]. Moreover, decomposition also reduces model complexity. For example, MAC-SQL [54] introduces a Decomposer agent to split the user’s query into subproblems, making sql generation for each part more manageable.
In general, the decomposition strategy divides the Text-to-SQL translation task into multiple subtasks, enabling each sub-module to focus on enhancing its specific output. However, this approach also raises computational costs, making model training and deployment more complex and resource-intensive.
@@
V-D Intermediate Representation for Text-to-SQL Translation
The Text-to-SQL is challenging due to the complexity and ambiguity of nl queries, coupled with the syntax-constrained nature of sql. To simplify this process, researchers have developed a grammar-free intermediate representation (IR) to bridge the “free-form” nl question and the “constrained and formal” sql. This IR provides a structured yet flexible format, capturing the essential components and relationships within an nl query without the strict syntax requirements of sql. Figure 8 shows two types of IR strategies, discussed below.
@@ V-D1 SQL-like syntax language
As shown in Figure 8(a), SQL-like syntax language is a simplified sql-like structure. Early approaches used information retrieval techniques to map the original question and schema data into this syntax [96, 114]. Subsequent research efforts have focused on consolidating or eliminating partial clauses or operations in sql queries to simplify SQL-like syntax language [98, 97]. SyntaxSQLNet [97] simplifies the syntax language by removing parts of the FROM and JOIN clauses. SemQL [98] removes the entire FROM, JOIN, ON, and GROUP BY clauses, and further merges the WHERE and HAVING conditions into a unified filtering representation. Recent research has focused on simplifying syntax languages to improve parsing efficiency [115]. NatSQL [100], a widely used SQL-like syntax language, eliminates uncommon sql operators and keywords, streamlining schema linking by minimizing necessary schema items. Combined with PLMs, NatSQL has achieved strong results on various benchmarks [7, 68].
@@ V-D2 SQL-like sketch structure
Early works applied fixed sketch rules and neural networks to map the nl into SQL-like sketch structure [116, 97]. SyntaxSQLNet [97] uses a syntax tree and a corresponding decoder, dividing the decoding into nine sub-modules that separately predict operators, keywords, and entities before combining them to generate the final sql. In recent years, the development of language models has allowed researchers to design more elaborate SQL-like sketch structures for parsing [1, 71, 44, 61]. CatSQL [71] constructs a more general template sketch with slots serving as initial placeholders. Its base model focuses on the parsing of nl to fill these placeholders, consequently decreasing the computational cost. Moreover, several recent works cover both SQL-like syntax language and SQL-like sketch transition methods. For instance, RESDSQL [7] introduces a rank-enhanced encoding and skeleton-aware decoding framework. During the decoding phase, its decoder initially generates the sql skeleton and then transforms it into the actual sql query. When combined with NatSQL, RESDSQL demonstrates the ability further to enhance the quality of sql query generation.
@@
VI Post-Processing Strategies for Text-to-SQL
After the Text-to-SQL model generates the sql, post-processing can refine it to better meet user expectations. This step involves leveraging additional information or models to enhance the sql, with a focus on sql correction, ensuring output consistency, and execution-guided checking.
Figure 9: Timeline for Text-to-SQL Benchmarks.